I need help on solving the linear system -4x+3y=1 3x+y=9 using substitution.
Answer by ad_alta(240)

There are several ways to do this, so I'll show one and let you figure out the others. Since -4x+3y=1, y=1/3+(4/3)x. We can substitute into the second equation and get 3x+1/3+(4/3)x=9. Thus, x=2 and y=3.